10. (RWJ, exercise 5.20) Allen Inc. is expected to pay an equal amount of dividends at the end of the first two years. Thereafter the dividend will grow at a constant rate of 4% indefinitely. The stock is currently traded at S30. What is the expected dividend per share for the next year if the required rate of return is 12%?
Answer #1

expected dividend pay (for 2nd year)

ke = d1 / p0

12% = d1 / 30

d1 = 30*.12 = 3.6

expected dividend pay (for 3rd year)

ke = d1 / p0 + g

12% = d1 / 30 + 4%

d1 = 30*.08 = 2.4

note:- in question not specified that dividend calculation required for which year. so we have calculated both year.